OpenStudy (anonymous):

(6y2-6/8y2+8y)/(3y-3/4y2+4y)

OpenStudy (camerondoherty):

\[ \Huge\frac{ \frac{ 6y^2-6 }{ 8y^2+8y } }{ \frac{ 3y-3 }{ 4y^2+4y } }\]Is it This?

OpenStudy (cp9454):

resolve in factors and then solve you will see every thin got cancelled only y+1 will be left

OpenStudy (cp9454):

\[\frac{ 6y^2-6 }{ 8y^2+8Y } * \frac{ 4y^2+4y }{ 3y-3 }\]

OpenStudy (cp9454):

this will go up as u see

OpenStudy (camerondoherty):

that is incorrect @cp9454 the answer would be \[\frac{ y^2 }{ y-1 }\]

OpenStudy (cp9454):

you need to check your answer lady @camerondoherty

OpenStudy (camerondoherty):

so do you

OpenStudy (camerondoherty):

its not multiplying its dividing

OpenStudy (camerondoherty):

you have the wrong equation

OpenStudy (cp9454):

\[\frac{ 6(y-1)(y+1) }{ 8y(y+1) } * \frac{ 4y(y+1) }{ 3(y-1) }\]

OpenStudy (cp9454):

\[\frac{ a }{ b \ \ \frac{ c }{ d } } = \frac{ a }{ b } * \frac{ d }{ c }\]

OpenStudy (cp9454):

it gets reversed and multiplied and the answer will not get effected.@camerondoherty

OpenStudy (camerondoherty):

what are yu tlking about? This is the equation\(\Huge\frac{ \frac{ 6y^2-6 }{ 8y^2+8y } }{ \frac{ 3y-3 }{ 4y^2+4y } }\)

OpenStudy (cp9454):

if you don't know don't argue with me. these are basic formulas. when you divide two fractions the fraction being divided gets reversed and multiplied.

OpenStudy (cp9454):

@camerondoherty

OpenStudy (cp9454):

how u got your answer @camerondoherty

OpenStudy (camerondoherty):

how did you get ur answer

OpenStudy (cp9454):

Well i have explained above, can't u see @camerondoherty

OpenStudy (cp9454):

how about we convert this in a challenge, if i won we will become friends and if you won i'll quit open study. @camerondoherty